You buy a bond with a par value of $1000 and a coupon rate of 9% with 21 coupons remaining. You hold the bond and receive 7 coupons. If the bond had a YTM of 8% when you bought it and 7% when you sold it, what was your annual holding period ROR?
Answer #1

Using financial calculator
Step 1: Purchase price
N=21
I/Y=8%
FV=-1000
PMT=-9%*1000
CPT PV=1100.17

Step 2: Selling price
N=14
I/Y=7%
FV=-1000
PMT=-9%*1000
CPT PV=1174.91

Step 3: ROR
N=7
PMT=9%*1000
PV=-1100.17
FV=1174.91
CPT I/Y=8.920767729%

Using Excel
=RATE(7,9%*1000,-PV(8%,21,-9%*1000,-1000),PV(7%,14,-9%*1000,-1000))
=8.920767729%